OpenStudy (anonymous):

Which expression represents (3xyw^4)(-6xy^2) OVER (-2xy^2)

OpenStudy (anonymous):

\[(\frac{\left( 3xyw^4 \right)\left( -6xy^2 \right)}{\left( -2xy^2 \right)}\]

OpenStudy (anonymous):

when you multiply (3xyw^4)(-6xy^2) you multiply the 3 and the -6, the x * x, the y * y and then you have the w^4, so what do we get there?

OpenStudy (anonymous):

sorry for the confusion, the reply is better to look at!

OpenStudy (anonymous):

it is

OpenStudy (mathstudent55):